A box contains $17.70 in nickels, dimes, and quarters. The number of dimes is 8 less than twice the number of nickels. The number of quarters is 2 more than the sum of the number of nickels and dimes. How many coins of each kind are there in the box?
Answer by Paul(988)

Numer of dimes = 2x-8
Nickels = x
Quarters = 2+(2x-8+x)
Total value x 100 = 1770
Nikel value x 100 = 5
Quarter value x 100 = 25
Dime value x 100 = 10
Equation:
5(x)+10(2x-8)+25(3x-6)=1770
5x+20x+75x-80-150=1770
100x=2000
x=20
2(20)-8=32
2+2(20)-8+20=54
Hence, there are 20 nickels, 54 quarters and 32 dimes.
Paul.